# Solving the Inverse Problem of Electrocardiography for Cardiac Digital Twins: A Survey

### v2

@@ 2.2.1 Spatial Regularization
To stabilize and improve the solution of the ECG inverse problem, spatial regularization methods are employed [57]. These methods introduce spatial constraints to promote smoothness in the estimated electrical activities across the heart surface at different orders of derivatives [58]. The commonly used spatial regularization methods include Tikhonov regularization [24], truncated singular value decomposition (TSVD) [59], truncated total least squares (TTLS) [60], L1-norm regularization [61], and total variation (TV) regularization [62]. These approaches aim to minimize the impact of noise and improve the accuracy of the inverse solution. For example, the Tikhonov regularization, including both zero-order and first-order variants, is the most widely used regularization technique [24, 63, 64, 65, 66, 67]. It minimizes the mean squared error from the body-heart transformation while simultaneously penalizing the L2 norm of the inverse HSP solution. The objective function can be formulated as,
@@
where the notation $\|\cdot\|^{2}$ represents the L2 norm, $\lambda_{Tikh}$ denotes the regularization coefficient, and $\Gamma$ is the operator constraining the HSP solution $u(\mathbf{s},t)$, which is an identity matrix for zero-order regularization and a spatial gradient operator for first-order regularization to increase the spatial smoothness of the inverse solution. This dual process effectively suppresses the unreliable components in the solution and enhances overall smoothness for improved accuracy. TSVD regularization method directly filters small singular values of the transfer matrix $\boldsymbol{H}$ and thus be robust to noises. TTLS regularization is effective in processing geometric errors that exist in the ECG inverse problem [60]. L1-norm and TV regularization techniques are useful for addressing specific challenges in the inverse problem, such as noise suppression, sparsity, and edge preservation.
@@ 2.2.2 Spatiotemporal Regularization
While spatial regularization methods are effective and widely used in ECG inverse problems, they often treat the problem as a static spatial problem and ignore the temporal dynamics of cardiac electrical activity. In some cases, this assumption may be reasonable, especially for certain types of cardiac abnormalities or when the temporal variations in the electrical activity are relatively slow compared to the sampling rate. However, in situations where the temporal dynamics are critical or when the electrical activity changes rapidly, temporal regularization methods that explicitly consider the time-varying nature of the problem may be more appropriate. These methods can help capture dynamic changes in electrical activity over time and provide more accurate and physiologically relevant solutions.
To fuse the space-time correlations into the inverse ECG problem, spatiotemporal regularizations are employed for robust estimation of HSP. For instance, besides employing Tikhonov regularization in the spatial domain, Messnarz et al. [68] introduced additional temporal constraints by assuming that cardiac electric potentials are non-decreasing during the depolarization phase. Yao et al. [69] proposed a spatiotemporal regularization method by regularizing both the spatial and temporal smoothness in the inverse ECG modelling, with the following objective function,
@@
where $\lambda_{s}$ and $\lambda_{t}$ are the spatial and temporal regularization terms, respectively, and $\omega$ denotes the selected time window to incorporate the temporal correlation. Splines-based methods can use spline functions to represent the electrical activity on the heart surface over time. They take advantage of the smoothness properties of splines to regularize the inverse solution, effectively capturing both spatial and temporal variations in the electrical potentials [70]. A Kalman filter reformulation can incorporate temporal information to regularize the inverse problem [71, 72]. Recently, to achieve regularization over space and time simultaneously, Cluitmans et al. [73] firstly applied regularization in the sparse wavelet domain for solving the ECG inverse problem, namely wavelet-promoted spatiotemporal regularization. Their experimental results indicate that the sparse representation of the cardiac source is beneficial for improving reconstruction accuracy in electrocardiographic imaging.
@@ 2.2.3 Model-based Regularization
Most existing regularization methods primarily rely on the optimization of the given observations, which are normally corrupted by noise in practice. In contrast, model-based regularization can integrate prior physiological knowledge about the electrical propagation inside the heart to constrain the solution space [74]. For instance, step jump functions and logistic functions have been employed to model the activation of action potential [75, 76]. Parameterized curves have been utilized to represent the wavefront velocity as trigonometric functions, while the potential has been modelled using the step response of a second-order linear system [72]. The ECG inverse problems can also be formulated via level-sets by evolving a boundary from an initial region to minimize a filtered residual error, allowing for iterative refinement of the solution. The major advantage of level-set-based methods is that no isotropy assumptions are required, allowing for more flexible modelling and accurate representation of complex geometries and phenomena within the heart [77]. To estimate transmural potentials throughout the myocardium, 3D biophysical EP simulation models have been employed to offer spatiotemporal constraints of the inverse issue [78, 79, 80].
While model-based regularization is effective in regularizing the ill-posed problem, their employed physiological models are developed for the general population [79, 74]. Specifically, the parameters of these models, which are typically linked to tissue characteristics (excitability and contractility of the heart muscle), are often assigned values commonly used in the literature. The general physiological models could overly constrain subject-specific information, leading to potential model errors and model-data mismatch [81]. For instance, the parameters controlling the configuration of the transmembrane potential (TMP) will differ throughout the heart, depending on whether the underlying tissue is in a healthy or diseased state. Therefore, Erem et al. [82] investigated how the solution changes when the assumption of uniform TMP amplitudes throughout the heart fails in the presence of infarction. Similarly, Xu et al. [81] and Dhamala et al. [83] emphasized the importance of uncertainty estimation in providing interpretable reconstructions of cardiac electrical activity.
@@ 2.2.4 Data-Driven Neural Networks
The notable rise in computational resources and the vast increase in available data have facilitated the development of data-driven models. Conventional deterministic models for ECG inverse problems typically rely on physics-based mathematical models, such as the bidomain, the monodomain, or the Eikonal model, to describe the propagation of electrical activity through the heart. In contrast, data-driven deterministic models can employ neural networks to directly learn the mapping from ECG data to the underlying electrical sources without relying on explicit physics-based equations. These models are highly flexible and can capture complex, non-linear relationships in the data. They do not rely on strong assumptions about the underlying physiology, making them suitable for cases where the true underlying model is not well-known, varies across patients, or displays unexpectedly abnormal behavior. Therefore, they have been employed to solve the ECG inverse problem. For example, Meister et al. [84] designed a deep learning method based on graph convolutional neural networks (CNNs) to estimate the depolarization patterns in the myocardium with scars. By leveraging the inherent graph structure of the cardiac anatomy, graph CNNs enable the extraction of spatial dependencies and patterns, facilitating a more nuanced representation of electrocardiographic data. Tenderini et al. [85] utilized a PDE-aware deep learning model for epicardial potential (EpiP) reconstruction from ECG data. Camara et al. [86] converted the BSP into an image represented by a 3D matrix and then employed CNNs to predict the locations of AF drivers from BSP. Neumann et al. [87] employed a reinforcement learning model to achieve multi-physics computational model personalization. Similarly, the time delay neural network can employ both current and previous timestep values of BSP to predict the HSP [88, 89, 90]. However, these purely data-driven models generally lack interpretability and may underperform in scenarios with limited data due to their data-intensive training requirements.
Figure 4: Example of the physic-informed neural networks for the reconstruction of activation time map. Here, the Eikonal equation has been employed as an additional physical constraint [91]. Image modified from Sahli [91] with permission.
Figure 5: Example of the physics-based model based on graph convolutional neural networks (CNNs) for the inverse inference of epicardial potential from body surface potential (BSP) [92]. Illustrations designed referring to Jiang et al. [92] and Bear et al. [93].
@@ 2.2.5 Physics-Informed Neural Networks
Physics-informed neural networks (PINNs) are data-efficient as they leverage both available data and the known physics equations [94]. They have recently been used to solve inverse problems with hidden physics, including the ill-posed ECG inverse problems [95]. This can be achieved either via introducing underlying physical constraints as an additional loss [95, 96], or designing specialized network architectures that satisfy the physics laws [92, 97]. For instance, Xie et al. [95] combined the data-driven loss and physics-based loss to inversely predict the heart-surface potentials from BSP and evaluated their framework in a 3D torso-heart geometry. Ye et al. [96] proposed a spatial-temporally adaptive PINN framework for solving both ECG forward and inverse problems over 3D bi-ventricular models. Sahli et al. [98] used PINNs in forward mode to estimate activation time maps (ATMs) and conduction velocity (CV) maps, as shown in Fig. 4. They incorporated a physics-informed regularization loss prescribed by the Eikonal equation, which describes the relationship between ATMs and the spatial gradient of CV. Similarly, Herrero et al. [97] developed a PINN model for action potential reconstruction and EP parameter estimation from sparse amounts of EP data. Although they did not inversely infer ATMs from BSP or ECG, a similar idea can be employed to solve the ECG inverse inference problem. Instead of introducing additional loss, Jiang et al. [92] proposed a spatial-temporal graph CNN-based non-Euclidean encoding-decoding model, where the geometry-dependent physics between BSP and HSP can be explicitly modelled via a bipartite graph over their graphical embeddings, as presented in Fig. 5. In general, the use of PINN in cardiac EP is mainly limited to 1D/ 2D in silico data, as it could be computationally expensive in realistic 2D/ 3D settings [97, 96].
@@ 2.3 Probabilistic Estimation Approaches
Unlike deterministic regularization techniques, probabilistic or stochastic estimation methods provide solutions in the form of probability distributions. Consequently, multiple potential solutions can be derived from the distributions, effectively addressing the inherent uncertainties within the data and the models relevant to the ECG inverse inference [19]. Specifically, the inverse problem can be formulated as the estimation of a Bayesian posterior probability distribution function (pdf), which can be defined as,
$p(\boldsymbol{u}|\boldsymbol{y})\propto p(\boldsymbol{y}|\boldsymbol{u},\epsilon)p(\boldsymbol{u}),$ (5)
where $p(\boldsymbol{u})$ is the prior, and the likelihood $p(\boldsymbol{y}|\boldsymbol{u},\epsilon)$ can be modeled as a normal distribution $\mathcal{N}(\boldsymbol{H}\boldsymbol{u},\epsilon\boldsymbol{I})$ [99]. One can predict the mode of the posterior distribution via maximum a posteriori (MAP) estimator, which is a point estimator that finds the parameter value that maximizes the posterior pdf given the observed data. When dealing with the complex dynamics of cardiac electrical activity, the Kalman filter serves as a specialized instance of the MAP estimator, enabling the integration of spatio-temporal information for the ECG inverse inference [100]. Alternatively, to estimate the entire posterior distribution, one can employ Markov Chain Monte Carlo sampling or variational inference for approximating analytically intractable posterior distributions. These approaches allow for a more comprehensive understanding of the cardiac electrical activity, taking into account the inherent variability and errors in the data and models, ultimately leading to more reliable and interpretable results.
@@ 2.3.1 Kalman Filter
The Kalman filter (KF) is a well-established tool for state estimation in dynamic systems. In the context of cardiac modelling, it implicitly estimates the posterior distribution of the underlying electrical activity given the observed ECG measurements. This is achieved by modelling the cardiac system as a high-dimensional stochastic state-space model, where the KF or its extensions can be effectively applied [79]. The key concept involves estimating the state $\boldsymbol{u}(\boldsymbol{s},t)$ by sequentially updating its estimate in the presence of noisy observations. The state transition is described as follows,
$\boldsymbol{u}(\boldsymbol{s},t)=\boldsymbol{F}\boldsymbol{u}(\boldsymbol{s},t-1)+\boldsymbol{w}_{t},$ (6)
where $\boldsymbol{F}$ refers to the state transition matrix (STM) and $\boldsymbol{w}_{t}$ is the process noise vector which models uncertainties and errors arising from the STM [101]. Then, the KF calculates the Kalman gain and uses it to update the state estimate based on observed data, considering spatio-temporal variations within cardiac source distributions. It provides not only a point estimate of the state but also a covariance matrix that represents the uncertainty associated with the estimate. By modelling both the dynamics of the system and the measurement noise probabilistically, KF inherently accounts for uncertainty in the estimation process.
To reduce computational complexity, researchers often focus on temporal relationships among spatially well-correlated regions, such as neighbouring regions or leads belonging to the same activation wavefront. They may also employ methods like expectation-maximization (EM) and residual-based techniques to estimate noise variances [101]. Other formulations, like the Duncan and Horn formulation of the KF used by Berrier et al. [71], focus on reconstructing specific electrograms. Wang et al. [79], in an attempt to integrate general prior knowledge with subject-specific data, used a statistical perspective to explicitly account for both model and data errors. They represented the cardiac electrophysiological system in a state-space form and conducted inverse inference of 3D TMP distributions via the unscented KF. A similar strategy was employed to define both the electrical and mechanical measurements as the state variables, which can be estimated via the unscented KF [102]. While the KF is a powerful tool for state estimation in cardiac modelling, it involves large matrices operation and may suffer from the initialization problem.

### v4

@@ 2.2.1 Spatial Regularization
To stabilize and improve the solution of the ECG inverse problem, spatial regularization methods are widely employed for ECGI [53]. These methods introduce spatial constraints to promote smoothness in the estimated electrical activities across the heart surface at different orders of derivatives [54]. The commonly used spatial regularization methods include Tikhonov regularization [25], truncated singular value decomposition (TSVD) [55], truncated total least squares (TTLS) [56], L1-norm regularization [57], and total variation (TV) regularization [58]. These approaches aim to minimize the impact of noise and improve the accuracy of the inverse solution. For example, the Tikhonov regularization, including both zero-order and first-order variants, is the most widely used regularization technique [25, 59, 60, 61, 50, 62]. It minimizes the mean squared error from the body-heart transformation while simultaneously penalizing the L2 norm of the inverse heart surface potential (HSP) solution. The objective function can be formulated as,
@@
where the notation $\|\cdot\|^{2}$ represents the L2 norm, $\lambda_{Tikh}$ denotes the regularization coefficient, and $\Gamma$ is the operator constraining the HSP solution $u(\mathbf{s},t)$, which is an identity matrix for zero-order regularization and a spatial gradient operator for first-order regularization to increase the spatial smoothness of the inverse solution. This dual process effectively suppresses the unreliable components in the solution and enhances overall smoothness for improved accuracy [63]. The choice of $\lambda_{\text{Tikh}}$ is crucial for enhancing reconstruction quality, and the L-curve method is commonly employed to determine the optimal value [64, 63]. TSVD regularization method directly filters small singular values of the transfer matrix $\boldsymbol{H}$ and thus be robust to noises. TTLS regularization is effective in processing geometric errors that exist in the ECG inverse problem [56]. L1-norm and TV regularization can tackle specific challenges in the inverse problems, such as noise suppression, sparsity, and edge preservation [65]. Note that although these methods have been extensively applied to HSP based solutions, they can be applied to other linear inverse problem representation, such as TMPs [66].
@@ 2.2.2 Spatiotemporal Regularization
While spatial regularization methods are effective and widely used in ECGI, they often overlook the temporal dynamics of cardiac electrical activity. To fuse the space-time correlations into the ECGI, spatiotemporal regularizations are therefore employed for robust estimation of cardiac sources. For instance, Twomey regularization was proposed as a modification of Tikhonov regularization to incorporates temporal a priori information [67, 68]. It adjusts the second term in Eq. 6 to minimize the difference between the solution and the prior estimate from the previous time instant. Alternative, one could introduce several additional constrains, including temporal constraints, to increase robustness to changes in the values of the regularization parameters [69, 70, 71]. Yao et al. [72] proposed a spatiotemporal regularization method by regularizing both the spatial and temporal smoothness, with the following objective function,
@@
where $\lambda_{s}$ and $\lambda_{t}$ are the spatial and temporal regularization terms, respectively, and $\omega$ denotes the selected time window to incorporate the temporal correlation. Recently, frequency-guided model have be widely applied for imaging the spatiotemporal behavior of the heart [48, 73, 74, 49]. Alternatively, splines-based methods can use spline functions to represent the electrical activity on the heart surface over time. They take advantage of the smoothness properties of splines to regularize the inverse solution, effectively capturing both spatial and temporal variations in the electrical potentials [75]. Erem et al. [76, 77] employed the temporal spline model to reconstruct TMP and ATM and localize pacing sites from BSP. However, their approach faced the challenge of determining the optimal number of spline functions in advance. To overcome this limitation, more flexible adaptive splines have been proposed to reconstruct cardiac sources, enabling data-driven selection of spline functions from a large set of basis elements through a stepwise process [75, 78]. Recently, to achieve regularization over space and time simultaneously, Cluitmans et al. [79] applied regularization in the sparse wavelet domain for reconstruct activation wavefronts and epicardial potential (EpiP), namely wavelet-promoted spatiotemporal regularization. Note that several studies have employed spatiotemporal Bayesian maximum a posteriori (MAP) or Kalman filters for ECGI [80, 81, 82], which however belong to probabilistic estimation and will be discussed further in Sec. 2.3.1 and 2.3.2.
Figure 3: Example of the physics-based model based on graph convolutional neural networks (CNNs) for the inverse inference of heart surface potential from body surface potential (BSP) [83]. Illustrations designed referring to Jiang et al. [83] and Bear et al. [84].
@@ 2.2.3 Model-based Regularization
Most existing regularization methods primarily rely on the optimization of the given observations, which are normally corrupted by noise in practice. In contrast, model-based regularization can integrate prior physiological knowledge about the electrical propagation inside the heart to constrain the solution space [85]. Compared to the above regularization methods that primarily focus on ECGI, model-based methods have been widely applied to both ECGI and EP parameter estimation.
For ECGI, model-based regularization can enhance the representation of cardiac electrical activity by incorporating detailed models of activation and wave propagation. For instance, step jump functions and logistic functions have been employed to model the activation of action potential [86, 87]. Parameterized curves have been utilized to represent the wavefront velocity as trigonometric functions, while the potential has been modelled using the step response of a second-order linear system [68]. The ECGI problems can also be formulated via level-sets by evolving a boundary from an initial region to minimize a filtered residual error, allowing for iterative refinement of the solution. The major advantage of level-set based methods is that no isotropy assumptions are required, allowing for more flexible modeling and accurate representation of complex geometries and phenomena within the heart [88, 89]. Cluitmans et al. [85] introduced a electrophysiology based regularization to model the cardiac surface potential using a set of basis vectors. The epicardial potential is then reconstructed as a sparse combination of these vectors, based on EP principles. To estimate transmural potentials throughout the myocardium, 3D biophysical EP simulation models have been employed to offer spatiotemporal constraints of the inverse issue [90, 91, 92]. Similar strategies have been previously employed for predicting onset activation location [93] and reconstrucing 3D activation time from BSP [94, 48, 95]. Furthermore, there exist a few model-based ECGI studies that optimize the model parameters via maximizing the correlation between simulated and measure BSP, which sometimes could be quite computational expensive [96, 97].
For EP parameter estimation, model-based regularization serves a different purpose. It aims to identify patient-specific parameters rather than applying generalized physiological models that are commonly used for ECGI [92]. This is generally achieved by optimization methods based on gradient descent [98, 99]. Furthermore, Chinchapatnam et al. [100] proposed an adaptive zonal decomposition iterative algorithm to estimate the local CV and conductivity within the Eikonal model. Similarly, Sermesant et al. [101] defined the regularization term as the energy of a simplified dynamical electromechanical model of the heart to estimate EASs and local CV.
@@ 2.2.4 Data-Driven Neural Networks
The notable increase in computational resources and available data have facilitated the development of data-driven models for the ECG inverse inference.
For ECGI, traditional methods often rely on regularization techniques and/ or biophysical models to stabilize the solution, but these approaches may struggle with accuracy and robustness in the presence of data noise or model errors. In contrast, data-driven neural networks can learn complex mappings between BSP and cardiac source directly from data, without the need for explicit biophysical modeling [102]. They do not rely on strong assumptions about the underlying physiology, making them suitable for cases where the true underlying model is not well-known, varies significantly across patients, or exhibits unexpected abnormal behavior. Therefore, they offer a promising alternative for solving the ECG inverse problem, particularly in scenarios with high variability or uncertainty in patient-specific cardiac physiology. The pioneering data-driven ECGI work was conducted by Yang et al. [103], who utilized the 12-lead ECG and convolutional neural networks (CNNs) to localize the origins of premature ventricular contractions (PVCs). They developed two distinct CNN models: one to classify the PVC origin across different segments and another to distinguish between endocardial (Endo) and epicardial (Epi) origins. Later, Meister et al. [104] designed a deep learning method based on graph CNNs to estimate the depolarization patterns in the myocardium with scars. Similary, Mu et al. [105] combined graph CNN and iterative soft threshold algorithm to reconstrcuct TMP and ATM from BSP. By leveraging the inherent graph structure of the cardiac anatomy, graph CNNs enable the extraction of spatial dependencies and patterns, facilitating a more nuanced representation of ECG data. Tenderini et al. [106] utilized a PDE-aware deep learning model for EpiP reconstruction from ECG data. Camara et al. [107] converted the BSP into an image represented by a 3D matrix and then employed CNNs to predict the locations of atrial fibrillation drivers from BSP. Cheng et al. [108] employed fast iterative shrinkage/thresholding network to reconstruct TMP from ECG. Chen et al. [109] tested the feasibility of several different neural networks, including CNN, fully connected neural network, and long short-term memory, for the reconstruction of EpiP and ATM as well as the localization of pacing sites. Furthermore, time delay neural networks have been employed to extract both current and previous timestep values of BSP to predict the HSP [110, 111, 112]. Interestingly, Karoui et al. [111] found that direct data-driven ATM reconstruction outperformed the two-stage approach, which first reconstructs HSP via a neural network and then estimates the ATM from the HSP using intrinsic deflection time. This is because small fluctuations in the reconstructed HSP can lead to errors in the subsequent computation of ATM.
For EP parameter estimation, the first deep learning based EP parameter personalization model was proposed by Neumann et al. [113], which however belongs to probabilistic estimation approaches (see Sec. 2.3). This is because they reformulated the EP parameter estimation problem into a Markov decision process and applied reinforcement learning to personalize the global CVs and Purkinje network [114]. Later, Giffard et al. [115] employed transfer learning to estimate local CVs and pacing sites from BSP for personalized cardiac resynchronization therapy (CRT). Nevertheless, these purely data-driven models generally lack interpretability and may underperform in scenarios with limited data due to their data-intensive training requirements.
@@ 2.2.5 Physics-Informed Neural Networks
Physics-informed neural networks (PINNs) are data-efficient as they leverage both available data and the known physics equations [116]. They have recently been used to solve inverse problems with hidden physics, including the ill-posed ECG inverse problems [117]. This can be achieved either via introducing underlying physical constraints as an additional loss [117, 118], or designing specialized network architectures that satisfy the physics laws [83, 119].
For ECGI, Xie et al. [117] combined the data-driven loss and physics-based loss to inversely predict the HSP from BSP and evaluated their framework in a 3D torso-heart geometry. Kashtanova et al. [120] proposed a APHYN-EP model, which decomposing the transmembrane dynamics into a physical term and a data-driven term, respectively. The data-driven deep learning component is tailored to capture information that the incomplete physical model fails to represent. This concept was later formally defined as the hybrid neural EP model by Jiang et al. [121]. Instead of introducing additional loss, Jiang et al. [83] proposed a spatial-temporal graph CNN-based model, where the geometry-dependent physics between BSP and HSP can be explicitly modelled via a bipartite graph over their graphical embeddings, as presented in Fig. 3.
For EP parameter estimation, Ye et al. [118] proposed a spatial-temporally adaptive PINN framework for both ECG simulation and AP model parameter estimation. Herrero et al. [119] developed a PINN model for action potential reconstruction and EP parameter estimation from sparse amounts of EP data. Instead, Sahli et al. [122] used PINNs in the forward model to interpolate activation time maps (ATMs) and estimate CV maps from a few sparse activation measurements. They incorporated a physics-informed regularization loss prescribed by the Eikonal equation, which describes the relationship between ATMs and the spatial gradient of CV. Although they did not inversely infer ATMs from BSP or ECG, a similar idea can be employed to achieve the ECG inverse inference. In general, PINNs for ECG inverse inference are mainly used for 1D/2D in silico data, as they face significant challenges, such as hyperparameter tuning, managing complex geometries, and spectral bias. These issues become more pronounced in realistic 2D/3D settings, due to their complex geometries and need for accurate spatial derivatives, leading to high computational requirements [119, 118].
@@ 2.3 Probabilistic Estimation Approaches
Unlike deterministic techniques, probabilistic estimation methods provide solutions in the form of probability distributions. Specifically, these approaches can naturally incorporate spatiotemporal information by using spatiotemporal priors or modeling state transitions [123]. In fact, deterministic methods such as Tikhonov regularization can be formulated as a special and simplified case of Bayesian MAP estimation under linear models with Gaussian priors [123]. In these cases, the regularization parameter effectively represents prior knowledge about the model. Probabilistic methods extend this by not only incorporating such prior information but also by providing a distribution of possible solutions. This distribution allows for comprehensive statistical analysis and helps address uncertainties inherent in the data and models relevant to ECG inverse inference [19].
2.3.1 Bayesian MAP Estimation
In Bayesian MAP estimation, the solution is selected by maximizing the posterior probability of the sources given the measurements. For ECGI, this maximization aims to find the most likely cardiac source distribution that matches both prior physiological knowledge and observed ECG measurements, given a specific forward model [124]. Greensite et al. [125] proposed using spatial and spatiotemporal Bayesian MAP methods for bioelectric (including ECG) inverse problems. Similarly, Zhou et al. [126] used a spatiotemporal Bayesian MAP approach to reconstruct EpiP and localize activation origins from BSP, which was later extended to the endocardial surface [127]. Instead of relying solely on BSP, Serinagaoglu et al. [128] utilized additional statistical prior information about the EpiP distribution, along with sparse EpiP measurements obtained from multielectrode coronary venous catheters, to enhance EpiP reconstruction within a Bayesian framework. Erenler et al. [123] assumed a time-invariant model when employing the Bayesian MAP for EGM reconstruction, which performed worse than Kalman filter solutions that consider the spatiotemporal variations of EGM.
@@ 2.3.2 Kalman Filter
The Kalman filter (KF) is a well-established method for estimating states in dynamic systems and has been used to improve ECGI by incorporating temporal information [129, 68, 123]. It implicitly estimates the electrical activity of the heart based on noisy ECG measurements by modeling the heart as a state-space system [92, 82]. The KF updates estimates of the state $\boldsymbol{u}(\boldsymbol{s},t)$ using the state transition matrix $\boldsymbol{F}$ and process noise $\boldsymbol{w}_{t}$ [130],
$\boldsymbol{u}(\boldsymbol{s},t)=\boldsymbol{F}\boldsymbol{u}(\boldsymbol{s},t-1)+\boldsymbol{w}_{t}.$ (8)
The filter calculates a point estimate and a covariance matrix representing the uncertainty. It effectively handles uncertainty by modeling both the system dynamics and measurement noise. To reduce computational complexity, researchers often focus on temporal relationships among spatially well-correlated regions, such as neighbouring regions or leads belonging to the same activation wavefront. They may also employ methods like expectation-maximization (EM) and residual-based techniques to estimate noise variances [130]. Other formulations, like the Duncan and Horn formulation of the KF used by Berrier et al. [129], focus on reconstructing specific EGMs. Wang et al. [92], in an attempt to integrate general prior knowledge with subject-specific data, used a statistical perspective to explicitly account for both model and data errors. They represented the cardiac EP system in a state-space form and conducted inverse inference of 3D TMP distributions via the unscented KF. A similar strategy was employed to define both the electrical and mechanical measurements as the state variables, which can be estimated via the unscented KF [131]. Huang et al. [132] proposed a data-driven Kalman filter model, consisting of a state transfer network and a Kalman gain network. Their method can consider the dynamic activation process of the cardiomyocyters when recovering TMP from BSP. While the KF is a powerful tool for state estimation in cardiac modeling, it involves large matrices operation and may suffer from the initialization problem.
For EP parameter estimation, to the best of our knowledge, only one study has used a reduced-order unscented KF to estimate cardiac tissue conductivity for personalizing a cardiac EP model [133]. In constrast, KF has been widely employed for the cardiac electromechanical model parameter estimation, such as contractility [134, 135].
Figure 4: Example of the MCMM application on the inverse inference of earliest activation sites (EASs) and CV from ECG. Here, MCMC has been used for the sampling of parameter sets of the population [1]. Image adapted from Camps et al. [1] with permission.